# IntelliProof: An Argumentation Network-based Conversational Helper for Organized Reflection

### v1

@@ Abstract
We present IntelliProof, an interactive system for analyzing argumentative essays through LLMs. IntelliProof structures an essay as an argumentation graph, where claims are represented as nodes, supporting evidence is attached as node properties, and edges encode supporting or attacking relations. Unlike existing automated essay scoring systems, IntelliProof emphasizes the user experience: each relation is initially classified and scored by an LLM, then visualized for enhanced understanding. The system provides justifications for classifications and produces quantitative measures for essay coherence. It enables rapid exploration of argumentative quality while retaining human oversight. In addition, IntelliProof provides a set of tools for a better understanding of an argumentative essay and its corresponding graph in natural language, bridging the gap between the structural semantics of argumentative essays and the user’s understanding of a given text. A live demo and the system are available here to try: https://intelliproof.vercel.app
@@
Intelliproof’s functionality spans argument creation, scoring, classification, and generation techniques. Each of the features elaborated on below is integrated within our GUI front-end.
@@
IntelliProof’s architecture consists of three core components. The frontend is built with Vite and React.js to create a dynamic user interface that handles all back-end API and database calls. The backend uses a Python server with FastAPI for handling requests and SupaBase (PostgreSQL) for managing user data such as profiles, evidence files, and graphs. For the large language model, we utilize GPT-4o via OpenAI’s Python library, chosen for its balance of performance, cost, and availability. The design is modular, allowing GPT-4o to be easily substituted with other locally or remotely deployed LLMs. [^1] We publicly release the source code for IntelliProof at https://github.com/collective-intelligence-lab/intelliproof

### v2

@@ Abstract
We present IntelliProof, an interactive system for analyzing argumentative essays through LLMs. IntelliProof structures an essay as an argumentation graph, where claims are represented as nodes, supporting evidence is attached as node properties, and edges encode supporting or attacking relations. Unlike existing automated essay scoring systems, IntelliProof emphasizes the user experience: each relation is initially classified and scored by an LLM, then visualized for enhanced understanding. The system provides justifications for classifications and produces quantitative measures for essay coherence. It enables rapid exploration of argumentative quality while retaining human oversight. In addition, IntelliProof provides a set of tools for a better understanding of an argumentative essay and its corresponding graph in natural language, bridging the gap between the structural semantics of argumentative essays and the user’s understanding of a given text.
Codegithub.com/collective-intelligence-lab/intelliproof
Demointelliproof.vercel.app
@@
Intelliproof’s functionality spans argument creation, scoring, classification, and generation techniques.
@@
IntelliProof’s architecture consists of three core components. The frontend is built with Vite and React.js to create a dynamic user interface that handles all back-end API and database calls. The backend uses a Python server with FastAPI for handling requests and SupaBase (PostgreSQL) for managing user data such as profiles, evidence files, and graphs. For the large language model, we utilize GPT-4o via OpenAI’s Python library, chosen for its balance of performance, cost, and availability. The design is modular, allowing GPT-4o to be easily substituted with other locally or remotely deployed LLMs.